In a labatory, a culture increases from 30 to 195 organisms in 5 hours. What is the hourly growth rate in the growth formula y=a(1+r)^t
Answer by nerdybill(7384)

In a labatory, a culture increases from 30 to 195 organisms in 5 hours. What is the hourly growth rate in the growth formula y=a(1+r)^t
.
y=a(1+r)^t
The problem gives us:
y as 195
a as 30
t as 5
.
195=30(1+r)^5
195/30 = (1+r)^5
6.5 = (1+r)^5
6.5^(1/5) = 1+r
6.5^(.2) = 1+r
6.5^(.2) - 1 = r
1.454 - 1 = r
0.454 = r
45.4% = r